If f is one-one onto and differentiable on R. then (f-1)(6)=1/f(6)
1
Expert's answer
2020-05-21T17:22:50-0400


False: Because f-1 here doesn't mean it is the reciprocal of the given function. Rather it is the inverse of the function which means you have to put the certain value in the inverse function you will get pre - image of the function .


Let's see the counter-example.

Suppose f(x)=x2, "x\geq 0,"

Then f-1(x)=√x

Now take the value x=6,we have

f-1(6)=√6 .....(1)

now f(6)=36

i.e. 1/f(6)=1/36 .....(2)

By equation (1) and (2),

f-1(6) is NOT equal to 1/f(6).

Answer: The given statement is false.